Three more than 4 times a number is equal to 9 less than twice the number. I s need to writh thisentance as an equation. Then I need to solve it.
Answer by mananth(16946)

let x be the number
4 times a number = 4x
3 more
=4x+3
............
twice the number = 2x
9 less = 2x-9
...
4x+3=2x-9
-2x
4x-2x+3=9
2x+3=9
-3
2x= 9-3
2x=6
/2
x=3 the number
m.ananth@hotmail.ca